OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution

OpenStudy (anonymous):

someone help please

OpenStudy (anonymous):

First subtract 15 from both sides, you get 3/x-8/ = 3 Now divide both sides by 3, you get /x-8/ = 1 Now set x - 8 = 1 and x - 8 = -1 So, x = 9 and x = 7 Seems like the minimum is 7 and maximum is 9.

OpenStudy (anonymous):

thank you so much im like far behind being far behind sucks badly